OpenStudy (anonymous):

solve for x tan^-1(x-1)/x-2 + tan^-1(x-1)/x+2 = pi

OpenStudy (anonymous):

i got x = 5/2;

OpenStudy (anonymous):

use the formulae - in denomenator we get = 1 - (x-1/x-2)*(x-1/x-2) tan pie = infinity so denominator shud be 0 solve and get 5/2 as x

OpenStudy (anonymous):

wat abt numerator

OpenStudy (anonymous):

no need they have given value of it as pie in the question

OpenStudy (anonymous):

no ...... u copied wrong quesstion i think

OpenStudy (anonymous):

u only wrote the question above

OpenStudy (anonymous):

oh its pi/4